a bookshelf has 84 books on it. There are twice as many non-fiction books as fiction books. How many fiction books are there on the bookshelf?

Answer by macston(5194)

F=Number of fiction books , 2F=no. of non-fiction books
total books =fiction + non-fiction=F+2F=84
3F=84 divide each side by 3
3F/3=84/3
F=28 ANSWER There are 28 fiction books on the shelf
CHECK:
There are 84 books with twice as many non-fiction as fiction
total=Fiction + non-fiction=28+2(28)=28+56=84